# A universal scheme to self-test any quantum state and measurement

### v1

@@ Abstract
The emergence of quantum devices has raised a significant issue: how to certify the quantum properties of a device without placing trust in it. To characterise quantum states and measurements in a device-independent way, up to some degree of freedom, we can make use of a technique known as self-testing. While schemes have been proposed to self-test all pure multipartite entangled states and real local rank-one projective measurements, little has been done to certify mixed entangled states, composite or non-projective measurements. By employing the framework of quantum networks, we propose a scheme that can be used to self-test any quantum state, projective measurement and rank-one extremal non-projective measurements. The quantum network considered in this work is the simple star network, which is implementable using current technologies. For our purposes, we also construct a family of Bell inequalities that can be used to self-test the two-dimensional tomographically complete set of measurements with an arbitrary number of parties.
@@
Introduction.—Demonstrating the strength of quantum systems over classical ones in tasks like computation and cryptography has been one of the most important advancements in quantum information theory. These quantum protocols are more efficient and secure when compared to the classical ones. To implement these protocols, it is thus necessary to certify the quantumness of a given device. The traditional methods of certifying a quantum state or a measurement rely on trusting the devices involved in the scheme, which are referred to as device-dependent schemes. For instance, any quantum state can be device-dependently verified if one has complete knowledge of the measurement apparatuses. However, from a practical perspective, it is unreasonable to completely trust the device. Therefore, it is highly desirable to devise certification schemes where the assumptions made on the devices are minimal, also referred to as device-independent (DI) certification schemes. DI schemes allow one to verify the non-trivial quantum properties of an unknown device by only observing the statistical data it generates. The essential resources for any DI scheme are Bell nonlocal correlations that cannot be reproduced via any local (or simply classical) hidden variable models [1, 2, 3]. For instance, observation of the violation of a Bell inequality is a DI certification of the presence of entanglement inside the device.
@@
Interestingly, the observation of Bell nonlocality in the data produced by a given device can provide a lot more information about the inner workings of the device than just the presence of entanglement. The strongest and most complete form of Bell nonlocality-based certification, known as self-testing [4, 5], allows for almost full characterization of the underlying quantum state and measurements performed on it. For this reason, in recent years, a plethora of schemes have been proposed to self-test pure quantum states or quantum measurements (see, e.g., Refs. [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20]). In particular, Refs. [13] and [19] propose schemes for self-testing any bipartite or multipartite pure entangled states, where the second method is based on the quantum networks scenario. The only scheme to self-test a mixed entangled state, in particular, a bound entangled state, was only recently proposed in Ref. [20] (see nevertheless Refs. [21, 22]). As for quantum measurements, apart from a bunch of schemes allowing to certify various examples of quantum measurements [16, 14, 23, 17, 24], including the composite ones [25, 26, 27, 20], a general method to certify any real local rank-one projective measurement has only very recently been designed in Ref. [24]. Despite this progress, there exist no single unified scheme that allows for certification of quantum measurements and quantum states (pure or mixed). Moreover, no general scheme for quantum networks-based certification of composite measurements has been proposed.
@@
The main aim of this work is to address the above challenges and provide a universal scheme that can be used to self-test any projective or rank-one extremal generalized quantum measurements (POVM) acting on arbitrary finite-dimensional Hilbert space as well as for any quantum state. For our purposes, we utilise the framework of quantum networks consisting of an arbitrary number of parties and sources (as depicted in Fig. 1 and explained later in the text). We first propose a Bell inequality that can be used to self-test the two-dimensional tomographically complete set of measurements. Based on this, we are then able to certify any projective or rank-one extremal generalized quantum measurement and eventually any quantum state. It is worth mentioning that, unlike the standard self-testing statements, our scheme requires additional causality constraints that are natural in quantum networks, in particular, that the sources are statistically independent. Our scheme is simple and can be implemented using the current infrastructure of quantum networks [28, 29, 30].
@@
$p(\mathbf{a}l|\mathbf{x}e)=\mathrm{Tr}\left(\rho_{AE}\bigotimes_{i=1}^{N}M_{a_{i}|x_{i}}\otimes R_{l|e}\right),$ (1)
@@
It is important to note that by using Eq. (1), these expectation values can be expressed as $\langle A_{1,x_{1}}\ldots A_{N,x_{N}}R_{l|e}\rangle=\mathrm{Tr}[(\bigotimes_{i=1}^{N}A_{i,x_{i}})\otimes R_{l|e}\rho_{AE}]$, where $A_{i,x_{i}}$ are quantum operators defined via the measurement elements as $A_{i,x_{i}}=M_{0|x_{i}}-M_{1|x_{i}}$ for every $x_{i}$ and $i$. Notice that in a particular case when $\{M_{0|x_{i}},M_{1|x_{i}}\}$ is a projective measurement, the corresponding $A_{x_{i}}$ becomes a standard hermitian observable. For our convenience, in the above representation Eve’s measurements remain represented in terms of the measurement elements $R_{l|e}$ instead of observables.
@@
If the above conditions (i) and (ii) are met one says that the reference state and measurements are self-tested in the actual experiment from the observed correlations. What is more, for any outcome $l$ of Eve’s measurement, the post-measurement states held by the external parties $A_{i}$,
$\tilde{\rho}^{l|e}_{A}=\frac{1}{\overline{P}(l|e)}\mathrm{Tr}_{E}\left[\left(\mathbbm{1}_{A}\otimes R_{l|e}\right)\bigotimes_{i=1}^{N}|\psi_{A_{i}E_{i}}\rangle\!\langle\psi_{A_{i}E_{i}}|\right],$ (5)
where $\overline{P}(l|e)$ is the probability that Eve observes the outcome $l$ after performing the $e-$th measurement satisfy
@@
where $\varrho_{A^{\prime\prime}}^{l|e}$ are some auxiliary states and $U_{A}=U_{A_{1}}\otimes\ldots\otimes U_{A_{N}}$ where $U_{A_{i}}$ are the same unitary transformations as those in Eqs. (3) and (4), then we say that the post-measurement states are self-tested too. It should be noted here that the above definition of self-testing the sources (3), measurements (4), and the post-measurement states (6) holds up to the complex conjugation of the reference states and measurements. The reason is that the correlations $\vec{p}$ are invariant under complex conjugation of quantum states and measurements.
@@
where $l_{1},l_{2},\ldots,l_{N}=0,1$ and $\overline{l}_{i}=1-l_{i}$ (for a proof see Fact 2 in [32]). Notice that, the measurements in Eq. (A universal scheme to self-test any quantum state and measurement) form a two-dimensional tomographically complete set of measurements for every party. We now comprise the result in the following theorem:
@@
Consider the scenario depicted in Fig. 1. If the Bell inequalities $\mathcal{I}_{l}$ (7) for any $l$ are maximally violated when Eve chooses the input $e=0$ with each outcome occurring with probability $\overline{P}(l|0)=1/2^{N}$, then the measurements of the external parties are equivalent, according to Eq. (4), to the reference measurements in (A universal scheme to self-test any quantum state and measurement), whereas the states $|\psi_{A_{i}E_{i}}\rangle$ are equivalent in the sense of Eq. (3) to the two-qubit maximally entangled state $|\phi^{+}\rangle=(|00\rangle+|11\rangle)/\sqrt{2}$.
@@
Sketch of the proof. Here we provide a sketch of the proof, deferring its full version to Appendix A. First, Eve performs the measurement corresponding to the input $e=0$ and observes an outcome $l$, in which way she prepares a post-measurement state shared by the external parties. Now, observation of maximal violation of the Bell inequalities $\mathcal{I}_{l}$ by the external parties allows one to self-test the post-measurement state as well as the external parties’ measurements. Then, the additional condition that the probabilities of obtaining the oucome $l$ by Eve when measuring $E_{0}$ is $1/2^{N}$ enables self-testing the states generated by the sources. This in turn also allows us to certify that Eve’s measurement $e=0$ is the Greenberger-Horne-Zeilinger (GHZ) basis $\{|\phi_{l}\rangle\!\langle\phi_{l}|\}$ where $|\phi_{l}\rangle$ are given in (10).
@@
Part 2. Now, that the measurements performed by the external parties as well as the states generated by the sources are self-tested, we can show how the quantum network can be used to self-test any projective quantum measurement or rank-one extremal POVM in the second measurement $E_{1}$ performed by the central party (see Fig. 2). To this end, let us first consider the case of projective measurements and introduce a $K$-outcome reference measurement $E^{\prime}_{1}=\{R^{\prime}_{l|1}\}$ defined on $\mathcal{H}_{N}=\left(\mathbbm{C}^{2}\right)^{\otimes N}$. While we assume this measurement to be projective, we do not impose any restrictions on the rank $r_{l}$ of the measurement elements $R_{l|1}^{\prime}$.
Now, in order to certify the reference measurement $E_{1}^{\prime}$ in the actual one $E_{1}=\{R_{l|1}\}$, we require that the assumptions of Theorem 1 are satisfied and additionally that for any outcome $l$, the following condition
$\sum_{i_{1},\ldots,i_{N}=0}^{3}f_{i_{1},\ldots,i_{N}}^{l}\left\langle\tilde{A}_{1,i_{1}}\otimes\bigotimes_{k=2}^{N}A_{k,i_{k}}\otimes(R_{l|1})_{E}\right\rangle=\frac{r_{l}}{2^{N}},$ (11)
is satisfied by the correlations $\vec{p}$ observed in the network, where $f_{i_{1},\ldots,i_{N}}^{l}$ are real coefficients that come from expanding $R^{\prime}_{l|1}$ in the $N$-fold tensor products of the qubit Pauli matrices [See Appendix B of [32] for details]. Moreover, $\tilde{A}_{1,0}$ and $\tilde{A}_{1,1}$ are defined in (8), whereas $\tilde{A}_{1,2}=A_{1,2}$ and $\tilde{A}_{1,3}=A_{k,3}=\mathbbm{1}$ for any $k$. Notice that the above conditions are met in the ideal situation in which the sources generate the two-qubit maximally entangled state, the external parties measure the observables (19), and, finally, Eve’s measurement $E_{1}$ is such that $R_{l|1}=(R^{\prime}_{l|1})^{*}$. Consequently, we find the following result.
Figure 2: Self-testing any quantum measurement. First, we have self-tested all the measurements performed by each external party $A_{i}$ as (A universal scheme to self-test any quantum state and measurement), as well as, the sources as the maximally entangled two-qubit state $|\phi_{A_{i}E_{i}}^{+}\rangle$. Thus, Eve’s first measurement is certified to the one in the GHZ basis (10). These are the ingredients sufficient to certify Eve’s arbitrary second measurement.
Theorem 2.
Consider the scenario depicted in Fig. 2 with states generated by the sources and measurements of the external parties being self-tested as in Theorem 1. If the observed correlations $\vec{p}$ additionally satisfy the conditions (11) for Eve’s measurement $E_{1}=\{R_{l|1}\}$, then this measurement is self-tested to be the reference projective measurement $E^{\prime}_{1}$, according to the definition (4).
The proof can be found in Appendix B of [32]. Importantly, the above statement can be extended to any projective measurement defined on a $D$-dimensional Hilbert space with arbitrary $D$. In fact, one can always embed the measurement $\{R^{\prime}_{l|1}\}$ into $\mathcal{H}_{N}=(\mathbbm{C}^{2})^{\otimes N}$ where $N$ is the smallest natural number such that $D<2^{N}$. Then, in order to make the measurement fully supported on $\mathcal{H}_{N}$ one completes it with an additional projective measurement element $M^{\perp}$ such that $\sum_{l}R^{\prime}_{l|1}+M^{\perp}=\mathbbm{1}_{2^{N}}$. This gives rise to an $(K+1)$-outcome projective measurement to which Theorem 2 applies.
Interestingly, in a similar way one can also certify an arbitrary extremal rank-one POVM’s. For this purpose, let us assume that Eve’s reference measurement $E_{1}^{\prime}=\{R^{\prime}_{l|1}\}$ is now some $K$-outcome extremal POVM where the measurement elements are given by $R^{\prime}_{l|1}=\lambda_{l}|\phi_{l}\rangle\!\langle\phi_{l}|$ for some $0\leq\lambda_{l}\leq 1$ and pure $N$-qubit pure states $|\phi_{l}\rangle$ such that $\sum_{l}\lambda_{l}|\phi_{l}\rangle\!\langle\phi_{l}|=\mathbbm{1}$.
Now, to self-test this reference measurement in $E_{1}$ the parties check whether for all outcomes $l=1,\ldots,K$ and sequences $i_{1},\ldots,i_{N}$ with $i_{j}=0,1,2,3$, the correlations $\vec{p}$ satisfy the following conditions,
@@
where $f^{l}_{i_{1},\ldots,i_{N}}$ are real numbers coming from the decompositions of the measurement operators $R^{\prime}_{l|1}$ in the $N$-fold tensor products of the qubit Pauli matrices. We can thus state our next result.
@@
It should be noticed that similarly to the previous case, a rank-one extremal POVM $E_{1}^{\prime}$ defined on $\mathbbm{C}^{D}$ for some $D$ can be embedded in $\mathcal{H}_{N}$ with $N$ being the minimal number such that $D<2^{N}$. To this end, one completes the measurement with additional $2^{N}-D$ rank-one projections $J_{i}$ so that the resulting measurement $\{R^{\prime}_{1|1},\ldots,R^{\prime}_{K|1},J_{1},\ldots,J_{2^{N}-D}\}$ is also extremal (see Appendix C for more details).
We thus obtain a general way of certifying any projective or rank-one extremal generalized measurement in the quantum networks scenario. Our scheme is based on violation of a class of simple Bell inequalities (7) and some additional conditions that the observed correlations $\vec{p}$ must satisfy: (11) for projective measurements or (12) for rank-one extremal POVMs. It is however worth noticing that the generality of our scheme comes at the price of the complexity that grows fast with the dimension of the measurement under consideration and thus with the number of involved external parties.
Part 3. Finally, we are going to show now that using the above setup one can self-test any quantum state. Consider again the network scenario depicted in Fig. 2. It is clear that Eve could remotely prepare different quantum states with the external parties by performing the concerned quantum measurement on the maximally entangled states generated by the sources. Thus, without loss of generality, the sources along with Eve can be considered a new preparation device. For instance, consider that Eve performs a projective measurement $E_{1}=\{|e_{1}\rangle\!\langle e_{1}|,\ldots,|e_{K}\rangle\!\langle e_{K}|\}$. Then the post-measurement state shared by the external parties when she obtains an outcome $l$ is $|e_{l}^{*}\rangle\!\langle e_{l}^{*}|$ where *** represents the complex conjugate. As a corollary to the previous statements, we can state the following fact.
@@
Consider the scenario depicted in Fig. 2. If the states generated by the sources along with the measurements performed by the external parties and Eve maximally violate the Bell inequalities (7) and satisfy the statistics (11) or (12) when Eve’s reference measurements are projective or rank-one extremal POVM respectively, then the post-measurement states with the external parties can be self-tested according to (6).
The proof of this statement can be found in Appendix D of [32]. It is important to emphasise here that, unlike the standard self-testing protocols, our scheme is not restricted to just pure entangled states. Rather, separable states and mixed states of any local dimension can also be self-tested with the external parties, according to (3).
@@
Importantly, however, the scheme is not restricted to pure states only and can straightforwardly extended to mixed states. Consider a reference mixed state $\rho=\sum_{k}p_{k}|\psi_{k}\rangle\!\langle\psi_{k}|$ acting on $\mathbbm{C}^{d}$ which we aim to self-test. For this purpose, Eve needs to perform an extremal $3d-$outcome POVM in the Hilbert space $\mathbbm{C}^{2d}$ on her share of the joint state $|\psi_{AE}\rangle$ to create $\rho$ at the external parties’ labs with the aid of post-processing. To construct this POVM, we first define two sets of $d$ mutually orthogonal vectors $\{|\psi_{k}\rangle\}$ and $\{|\phi_{k}\rangle\}$ such that $\langle\psi_{k}|\phi_{k^{\prime}}\rangle=0$ for any $k,k^{\prime}$. One then considers a pair $\{|\psi_{k}\rangle,|\phi_{k}\rangle\}$ to construct the a three-outcome trine POVM [cf. Ref. [33]] as

### v2

# A universal scheme to self-test any quantum state and extremal measurement

@@ Abstract
The emergence of quantum devices has raised a significant issue: how to certify the quantum properties of a device without placing trust in it. To characterise quantum states and measurements in a device-independent way, up to some degree of freedom, we can make use of a technique known as self-testing. While schemes have been proposed to self-test all pure multipartite entangled states (up to complex conjugation) and real local rank-one projective measurements, little has been done to certify mixed entangled states, composite or non-projective measurements. By employing the framework of quantum networks, we propose a scheme for self-testing (up to complex conjugation) arbitrary extremal measurements, including the projective ones, but also in an indirect way any quantum states, including the mixed ones. The quantum network considered in this work is the simple star network, which is implementable using current technologies. For our purposes, we also construct a scheme that can be used to self-test the two-dimensional tomographically complete set of measurements with an arbitrary number of parties.
@@
Introduction.—Demonstrating the strength of quantum systems over classical ones in tasks like computation and cryptography has been one of the most important advancements in quantum information theory. These quantum protocols are more efficient and secure when compared to the classical ones. To implement these protocols, it is thus necessary to certify the quantumness of a given device. The traditional methods of certifying a quantum state or a measurement rely on trusting the devices involved in the scheme, which are referred to as device-dependent schemes. For instance, any quantum state can be device-dependently verified if one has complete knowledge of the measurement apparatuses. However, from a practical perspective, it is unreasonable to completely trust the device. Therefore, it is highly desirable to devise certification schemes where the assumptions made on the devices are minimal, also referred to as device-independent (DI) certification schemes. DI schemes allow one to verify the non-trivial quantum properties of an unknown device by only observing the statistical data it generates. The essential resources for any DI scheme are Bell nonlocal correlations that cannot be reproduced via any local (or simply classical) hidden variable models [1, 2, 3]. For instance, observation of the violation of a Bell inequality certifies in a DI manner that the device operates on an entangled state.
Interestingly, the observation of Bell nonlocality in the data produced by a given device can provide a lot more information about its internal working than just the presence of entanglement. The strongest and most complete form of Bell nonlocality-based certification, known as self-testing [4, 5], allows for almost full characterization of the underlying quantum state and measurements performed on it.
@@
In recent years, a plethora of schemes have been proposed to self-test pure quantum states or quantum measurements (see, e.g., Refs. [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20]). In particular, Refs. [13] and [19] propose self-testing strategies for any bipartite or multipartite pure entangled states, where the second method is based on the quantum networks scenario. The only scheme to self-test a mixed entangled state, in particular, a bound entangled state, was only recently proposed in Ref. [20] (see nevertheless Refs. [21, 22]). As for quantum measurements, apart from a bunch of schemes allowing to certify various examples of quantum measurements [16, 14, 23, 17, 24], including the composite ones [25, 26, 27, 20], a general method to certify any real local rank-one projective measurement has only very recently been designed in Ref. [24]. Despite this progress, there exist no single unified scheme that allows for certification of quantum measurements and quantum states (pure or mixed). Moreover, no general scheme for quantum networks-based certification of composite measurements has been proposed.
@@
The main aim of this work is to address the above challenges and provide a universal scheme that can be used to self-test (up to complex conjugate) any extremal generalized quantum measurement (POVM) acting on arbitrary finite-dimensional Hilbert space Our scheme also includes all projective measurements as any projective measurement in extremal. Interestingly, our approach allows also for an indirect certification of any, even mixed quantum state. For our purposes, we utilise the framework of quantum networks consisting of an arbitrary number of parties and sources (as depicted in Fig. 1 and explained later in the text). We first propose a Bell inequality that can be used to self-test the two-dimensional tomographically complete set of measurements with arbitrary number of parties. Based on this, we are then able to certify any extremal POVM and eventually any quantum state. It is worth mentioning that, unlike the standard self-testing statements, our scheme requires additional causality constraints that are natural in quantum networks, in particular, that the sources are statistically independent. Our scheme is simple and can be implemented using the current infrastructure of quantum networks [28, 29, 30].
@@
$p(\mathbf{a}l|\mathbf{x}e)=\mathrm{Tr}\left[\rho_{AE}\left(\bigotimes_{i=1}^{N}M_{a_{i}|x_{i}}\otimes R_{l|e}\right)\right],$ (1)
@@
It is important to note that by using Eq. (1), these expectation values can be expressed as $\langle A_{1,x_{1}}\ldots A_{N,x_{N}}R_{l|e}\rangle=\mathrm{Tr}[(\bigotimes_{i=1}^{N}A_{i,x_{i}})\otimes R_{l|e}\rho_{AE}]$, where $A_{i,x_{i}}$ are quantum operators defined via the measurement elements as $A_{i,x_{i}}=M_{0|x_{i}}-M_{1|x_{i}}$ for every $x_{i}$ and $i$. Notice that in a particular case when the measurement $\{M_{0|x_{i}},M_{1|x_{i}}\}$ is projective, the corresponding operator $A_{x_{i}}$ becomes a standard hermitian observable with ±$\pm 1$ eigenvalues. For our convenience, in the above representation Eve’s measurements remain represented in terms of the measurement elements $R_{l|e}$ instead of observables.
@@
If the above conditions (i) and (ii) are met one says that the reference state and measurements are self-tested in the actual experiment from the observed correlations. What is more, for any outcome $l$ of Eve’s measurement, the post-measurement states $\rho^{l|e}_{A}$ held by the external parties $A_{i}$, satisfy
@@
where $\tilde{\rho}^{l|e}_{A^{\prime}}$ are the reference post-measured states and $\varrho_{A^{\prime\prime}}^{l|e}$ are some auxiliary states and $U_{A}=U_{A_{1}}\otimes\ldots\otimes U_{A_{N}}$ where $U_{A_{i}}$ are the same unitary transformations as those in Eqs. (3) and (4), then we say that the post-measurement states are self-tested too. It should be noted here that the above definition of self-testing the sources (3), measurements (4), and the post-measurement states (5) holds up to the complex conjugation of the reference states and measurements. The reason is that the correlations $\vec{p}$ are invariant under complex conjugation of quantum states and measurements.
@@
where $l_{1},l_{2},\ldots,l_{N}=0,1$ and $\overline{l}_{i}=1-l_{i}$ (for a proof see Fact 2 in [32]). Notice that, the measurements in Eq. (A universal scheme to self-test any quantum state and extremal measurement) form a two-dimensional tomographically complete set of measurements for every party. We now comprise the result in the following theorem:
@@
Consider the scenario depicted in Fig. 1. If the Bell inequalities $\mathcal{I}_{l}$ (6) for any $l$ are maximally violated when Eve chooses the input $e=0$ with each outcome occurring with probability $\overline{P}(l|0)=1/2^{N}$, then the measurements of the external parties are equivalent, according to Eq. (4), to the reference measurements in (A universal scheme to self-test any quantum state and extremal measurement), whereas the states $|\psi_{A_{i}E_{i}}\rangle$ are equivalent in the sense of Eq. (3) to the two-qubit maximally entangled state $|\phi^{+}\rangle=(|00\rangle+|11\rangle)/\sqrt{2}$.
@@
Sketch of the proof. Here we provide a sketch of the proof, deferring its full version to Appendix A. First, Eve performs the measurement corresponding to the input $e=0$ and observes an outcome $l$, in which way she prepares a post-measurement state shared by the external parties. Now, observation of maximal violation of the Bell inequalities $\mathcal{I}_{l}$ by the external parties allows one to self-test the post-measurement state as well as the external parties’ measurements. Then, the additional condition that the probabilities of obtaining the oucome $l$ by Eve when measuring $E_{0}=\{R_{l|0}\}$ is $1/2^{N}$ enables self-testing the states generated by the sources. This in turn also allows us to certify that Eve’s measurement $E_{0}$ is the Greenberger-Horne-Zeilinger (GHZ) basis $\{|\phi_{l}\rangle\!\langle\phi_{l}|\}$ where $|\phi_{l}\rangle$ are given in (9).
@@
Part 2. Now, that the measurements performed by the external parties as well as the states generated by the sources are self-tested, we can show how the quantum network can be used to self-test any extremal POVM in the second measurement $E_{1}$ performed by the central party (see Fig. 2). For this purpose, let us assume that Eve’s reference measurement $E_{1}^{\prime}=\{R^{\prime}_{l|1}\}$ is now some $K$-outcome extremal POVM defined on $\mathcal{H}_{N}=\left(\mathbbm{C}^{2}\right)^{\otimes N}$. Now, to self-test this reference measurement in $E_{1}$ the parties check whether for all outcomes $l=1,\ldots,K$ and sequences $i_{1},\ldots,i_{N}$ with $i_{j}=0,1,2,3$, the correlations $\vec{p}$ satisfy the following conditions,
@@
where $f^{l}_{i_{1},\ldots,i_{N}}$ are real numbers coming from the decompositions of the measurement operators $R^{\prime}_{l|1}$ in the $N$-fold tensor products of the qubit Pauli matrices [see Appendix B of [32] for details]. We can thus state our next result.
@@
Figure 2: Self-testing any quantum measurement. First, we have self-tested all the measurements performed by each external party $A_{i}$ as (A universal scheme to self-test any quantum state and extremal measurement), as well as, the sources as the maximally entangled two-qubit state $|\phi_{A_{i}E_{i}}^{+}\rangle$. Thus, Eve’s first measurement is certified to the one in the GHZ basis (9). These are the ingredients sufficient to certify Eve’s arbitrary second measurement.
The proof of this statement can be found in Appendix B of [32]. Importantly, the above statement can be extended to any extremal measurement defined on a $D$-dimensional Hilbert space with arbitrary $D$. In fact, one can always embed the measurement $\{R^{\prime}_{l|1}\}$ into $\mathcal{H}_{N}=(\mathbbm{C}^{2})^{\otimes N}$ where $N$ is the smallest natural number such that $D<2^{N}$. Then, in order to make the measurement fully supported on $\mathcal{H}_{N}$ one completes it with an additional projector $M^{\perp}$ such that $\sum_{l}R^{\prime}_{l|1}+M^{\perp}=\mathbbm{1}_{2^{N}}$. This gives rise to an $(K+1)$-outcome extremal measurement to which Theorem 2 applies. Note that, in the case of composite measurements acting on $\bigotimes_{i}\mathbbm{C}^{D_{i}}$ each of the $i-$th local Hilbert spaces need to be embedded in $(\mathbbm{C}^{2})^{\otimes n_{i}}$ such that $\sum_{i}n_{i}=N$.
We thus obtain a general way of certifying any extremal generalized measurement in the quantum networks scenario. Our scheme is based on violation of a class of simple Bell inequalities (6) and some additional conditions (10) that the observed correlations $\vec{p}$ must satisfy. It is however worth noticing that the generality of our scheme comes at the price of the complexity that grows fast with the dimension of the measurement under consideration and thus with the number of involved external parties.
Part 3. Finally, we are going to show now that using the above setup one can self-test any quantum state. Consider again the network scenario depicted in Fig. 2. It is clear that Eve could remotely prepare different quantum states with the external parties by performing the concerned quantum measurement on the maximally entangled states generated by the sources. Thus, without loss of generality, the sources along with Eve can be considered a new preparation device. For instance, consider that Eve performs a projective measurement $E_{1}=\{|e_{1}\rangle\!\langle e_{1}|,\ldots,|e_{K}\rangle\!\langle e_{K}|\}$. Then the post-measurement state shared by the external parties when she obtains an outcome $l$ is $|e_{l}^{*}\rangle\!\langle e_{l}^{*}|$ where * represents the complex conjugate. As a corollary to the previous statements, we can state the following fact.
@@
Consider the scenario depicted in Fig. 2. If the states generated by the sources along with the measurements performed by the external parties and Eve maximally violate the Bell inequalities (6) and satisfy the statistics (10) when Eve’s reference measurements are extremal, then the post-measurement states with the external parties can be self-tested according to (5).
The proof of this statement can be found in Appendix C of [32]. It is important to emphasise here that, unlike the standard self-testing protocols, our scheme is not restricted to just pure entangled states. Rather, separable states and mixed states of any local dimension can also be self-tested with the external parties, according to (3).
@@
Importantly, however, the scheme is not restricted to pure states only and can straightforwardly extended to mixed states. Consider a reference mixed state $\rho=\sum_{k}p_{k}|\psi_{k}\rangle\!\langle\psi_{k}|$ acting on $\mathbbm{C}^{d}$ which we aim to self-test. For this purpose, Eve needs to perform an extremal $3d-$outcome POVM in the Hilbert space $\mathbbm{C}^{2d}$ on her share of the joint state $|\psi_{AE}\rangle$ to create $\rho$ at the external parties’ labs with the aid of post-processing. To construct this POVM, we first define two sets of $d$ mutually orthogonal vectors $\{|\psi_{k}\rangle\}$ and $\{|\phi_{k}\rangle\}$ such that $\langle\psi_{k}|\phi_{k^{\prime}}\rangle=0$ for any $k,k^{\prime}$. One then considers a pair $\{|\psi_{k}\rangle,|\phi_{k}\rangle\}$ to construct a three-outcome trine POVM [cf. Ref. [33]] as